# Xingjian Wu, Jianxin Jin, Wanghui Qiu, Peng Chen, Yang Shu, Bin Yang, Chenjuan Guo

### v1

@@ 2.2 Time Series Foundation Models
To support cross-domain generalization, unimodal Time Series Foundation Models are widely studied. The majority of them adopt Tranformer-based architectures, which are pretrained on time series corpus of billion- or trillion- scale to obtain the strong generalization capabilities. Among them, Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), UniTS (Gao et al., 2024), Timer (Liu et al., 2024d), and TimesFM (Das et al., 2024) demonstrate strong zero-shot forecasting performance on unimodal tasks, even outperforming those full-shot supervised models in many cases. Considering the forecasting paradigm, Sundial, MOIRAI, Chronos, and Lag-Llama (Rasul et al., 2023) also support probabilistic forecasting, which provides additional robustness and versatility for decision-making. Despite their endeavors to enhance cross-domain generalization capabilities, when historical series exhibit similarities, the forecasts they generate remain static. This lack of adaptability renders them unable to accommodate the diverse and changing real-world domains.
@@ 4.1 Experimental Settings
Baselines. We compare Aurora with 10 well-known unimodal time series foundation models, including Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TimesFM (Das et al., 2024), Timer (Liu et al., 2024d), UniTS (Gao et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023). We also consider 8 strong end-to-end supervised models, including multimodal ones like GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), and Time-VLM (Zhong et al., ), and unimodal ones like TSDiff (Kollovieh et al., 2023), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and GRU MAF (Rasul et al., 2021b). The detailed information is provided in Appendix A.
@@ A.2 Baselines
For zero-shot forecasting, we compare Aurora against 10 advanced foundation models: Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Timer (Liu et al., 2024d), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), TimesFM (Das et al., 2024), Time-MoE (Shi et al., 2024b), UniTS (Gao et al., 2024), and Lag-Llama (Rasul et al., 2023). In probabilistic forecasting and multimodal forecasting tasks, we also compare Aurora with 8 End-to-end supervised models: GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), Time-VLM (Zhong et al., ), TSDiff (Kollovieh et al., 2023), TimeGrad (Rasul et al., 2021a), CSDI (Tashiro et al., 2021), and GRU MAF (Rasul et al., 2021b). The corresponding codebases and implementation details are summarized in Table 5.
@@ Appendix C Related Works
Time series forecasting (TSF) predicts future observations based on historical observations. Studies have shown that learned features may perform better than human-designed features (Qiu et al., 2025b; a; Liu et al., 2025a; Lu et al., 2023; 2024a; 2024b; Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025; Li et al., 2025b; Wang et al., 2025a; Wu et al., 2025c; a; 2024; Qiu et al., 2025c; Wu et al., 2025b; Qiu et al., 2025e; Dai et al., 2024c; Hu et al., 2025a; Dai et al., 2024d; Liu et al., 2025c; Sun et al., 2025b; a; Niu et al., 2025). Leveraging the representation learning of deep neural networks (DNNs), many deep learning-based methods emerge. TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, using CNNs or RNNs to capture temporal dependencies. Additionally, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025c), LightGTS (Wang et al., 2025b), and MagicScaler (Pan et al., 2023) capture complex relationships between time points more accurately, significantly improving forecasting performance. MLP-based methods, including DUET (Qiu et al., 2025d), AMD (Hu et al., 2025a), SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt simpler architectures with fewer parameters but still achieve highly competitive forecasting accuracy.

### v2

@@ 2.2 Time Series Foundation Models
To support cross-domain generalization, unimodal Time Series Foundation Models are widely studied. The majority of them adopt Tranformer-based architectures, which are pretrained on time series corpus of billion- or trillion- scale to obtain the strong generalization capabilities. Among them, Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), UniTS (Gao et al., 2024), Timer (Liu et al., 2024e), and TimesFM (Das et al., 2024) demonstrate strong zero-shot forecasting performance on unimodal tasks, even outperforming those full-shot supervised models in many cases. Considering the forecasting paradigm, Sundial, MOIRAI, Chronos, and Lag-Llama (Rasul et al., 2023) also support probabilistic forecasting, which provides additional robustness and versatility for decision-making. Despite their endeavors to enhance cross-domain generalization capabilities, when historical series exhibit similarities, the forecasts they generate remain static. This lack of adaptability renders them unable to accommodate the diverse and changing real-world domains.
@@ 4.1 Experimental Settings
Baselines. We compare Aurora with 10 well-known unimodal time series foundation models, including Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Time-MoE (Shi et al., 2024b), MOIRAI (Woo et al., 2024), TimesFM (Das et al., 2024), Timer (Liu et al., 2024e), UniTS (Gao et al., 2024), Chronos (Ansari et al., ), and Lag-Llama (Rasul et al., 2023). We also consider 8 strong end-to-end supervised models, including multimodal ones like GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), and Time-VLM (Zhong et al., ), and unimodal ones like TSDiff (Kollovieh et al., 2023), CSDI (Tashiro et al., 2021), TimeGrad (Rasul et al., 2021a), and GRU MAF (Rasul et al., 2021b). The detailed information is provided in Appendix A.
@@ A.2 Baselines
For zero-shot forecasting, we compare Aurora against 10 advanced foundation models: Sundial (Liu et al., 2025d), VisionTS (Chen et al., 2024a), ROSE (Wang et al., 2025c), Timer (Liu et al., 2024e), MOIRAI (Woo et al., 2024), Chronos (Ansari et al., ), TimesFM (Das et al., 2024), Time-MoE (Shi et al., 2024b), UniTS (Gao et al., 2024), and Lag-Llama (Rasul et al., 2023). In probabilistic forecasting and multimodal forecasting tasks, we also compare Aurora with 8 End-to-end supervised models: GPT4MTS (Jia et al., 2024), TATS (Li et al., 2025d), CALF (Liu et al., 2025a), Time-VLM (Zhong et al., ), TSDiff (Kollovieh et al., 2023), TimeGrad (Rasul et al., 2021a), CSDI (Tashiro et al., 2021), and GRU MAF (Rasul et al., 2021b). The corresponding codebases and implementation details are summarized in Table 5.
@@ Appendix C Related Works
Time Series Analysis assumes a position of utmost significance within a wide array of fields, including the economy (Qiu et al., 2025b; a; Liu et al., 2025a), transportation (Wu et al., 2025c; 2024), health (Lu et al., 2023; 2024a; 2024b), weather (Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses numerous crucial tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024d), anomaly detection (Wang et al., 2025a; Qiu et al., 2025c; Wu et al., 2025b), classification (Liu et al., 2024d), imputation (Wu et al., 2023), and others (Qiu et al., 2025e; Wu et al., 2025a). Among these tasks, Time Series Forecasting is the one most widely applied in real-world situations.
Time series forecasting (TSF) entails the prediction of future observations based on historical ones. Research findings have indicated that features learned through certain methods may exhibit superior performance compared to human-designed features (Dai et al., 2024c; Hu et al., 2025a; Dai et al., 2024d; Liu et al., 2025c; Sun et al., 2025b; a; Niu et al., 2025). By harnessing the representation learning capabilities of deep neural networks (DNNs), a multitude of deep-learning based approaches have emerged. For instance, TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) model time series as vector sequences, employing CNNs or RNNs to capture temporal dependencies. Moreover, Transformer architectures, including Informer (Zhou et al., 2021), Dsformer (Yu et al., 2023), TimeFilter (Hu et al., 2025b), TimeBridge (Liu et al., 2025b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), PatchTST (Nie et al., 2023), ROSE (Wang et al., 2025c), LightGTS (Wang et al., 2025b), and MagicScaler (Pan et al., 2023), are capable of more accurately capturing the complex relationships between time points, thereby significantly enhancing forecasting performance. MLP-based methods, such as DUET (Qiu et al., 2025d), AMD (Hu et al., 2025a), SparseTSF (Lin et al., 2024a), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters yet still manage to achieve highly competitive forecasting accuracy.